[Hallc_running] [New Logentry] RC update

simona at jlab.org simona at jlab.org
Tue Jan 25 08:55:02 EST 2022

Logentry Text:
Yesterday we ran well until ~10:30 am when beam delivery to Hall C was interrupted for the Hall B pass change. We got beam back 4 hours later and resumed production on LH2 at Q2=5, SHMS left. During the second half of the swing shift and throughout owl beam delivery to Hall C was not that great due to RF instabilities. We got about 5.5 hours of ABUs during swing and just 2.2 hours during owl. There were no downs on the hall side. We got 57% of the goal statistics on LH2 so far.

The 2L18 cryo bump exercise started this morning. We will be going to controlled access as soon as Radcon can support the survey of the hall this morning. Please do not access the hall without informing the RC first.

Beam delivery to the halls will resume tomorrow during day shift.

The Hall C counting house will be staffed with TOs at all times to monitor the target. The SLs for today day and swing and tomorrow owl are on standby. 

We will have a brief RC meeting today at 10 am: https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcEE0QT09


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3976948
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220125/89ecc096/attachment.html>

More information about the Hallc_running mailing list